OpenStudy (anonymous):

Describe how to prepare 1.00 L of 0.646 M HCl solution, starting with 2.00 M HCl solution.

OpenStudy (photon336):

It's been a while but I think it may be this: \[M_{1}V_{1} = M_{2}V_{2}\] \[\frac{ M_{1}V_{1} }{ M_{2} } = V_{2}\] \[\frac{ 1.0L*0.646M }{ 2.00M } = 0.323 L \] we would need 323 mL of a 2.0M solution @Rushwr your help is needed @Shikamaru11 my good friend thoughts?

OpenStudy (shikamaru11):

well @Photon336 Sir this rite no doubt :)

OpenStudy (rushwr):

@Photon336 Absolutely correct :) Dude u are awesome at chemistry.

OpenStudy (anonymous):

Hmm...

OpenStudy (joannablackwelder):

Yep, diluted with water to 1 L